# TIPO: Text to Image with Text Presampling for Prompt Optimization

### v3

@@ F.5 Statistical Significance
F.6 Survey Response Examples
@@ F.7 Conclusion
Appendix G Ablation Study on TIPO
@@
G.1 Experimental Setup
Prompt Variants.
@@ G.1 Experimental Setup
Inference Procedure.
Prompts are fed into our image-generation pipeline under identical model settings (classifier free guidance, sampler, steps, etc.), using the v-parameterized variant of Illustrious v3.5[49]. We focus on how TIPO modifications alter the generation outcomes and whether they introduce additional computational overhead.
G.2 Evaluation Metrics
@@
G.3 Results & Discussion
Impact on Aesthetics.
@@ G.3 Results & Discussion
Improvements in Corruption Score.
@@
G.4 Speed Test and Overhead Analysis
@@
Memory Footprint.
We also confirm that TIPO’s overhead in terms of VRAM usage is minimal. The practical adoption of TIPO in pipelines has shown no critical memory concerns, which aligns with our measurements.
G.5 Conclusion of Ablation
@@
Appendix H Topic Distribution Visualization
Latent Dirichlet Allocation (LDA) [9] is a generative probabilistic model for topic modeling [32], which assumes that each document is a mixture of topics, with each topic represented by a distribution over words. LDA uncovers hidden thematic structures by analyzing word co-occurrence patterns, while methods like TF-IDF and TextRank [43] enhance its ability to extract meaningful insights from large textual datasets. We implemented a multi-stage topic modeling and clustering methodology using LDA to extract varying numbers of topics (20, 30, 50, and 100) from the corpus. This approach focuses on identifying significant representative words while filtering out stop words and irrelevant terms to ensure meaningful topic classification.
We empirically assessed whether the resulting topics were sufficiently large and diverse by employing multi-level topic analysis. This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics [62].
To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering [70]. We grouped the initially extracted topics into five major clusters using k-means clustering. We evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [27], shown in Table 10, 11, and 12. Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset. This two-tiered approach allows for a more nuanced analysis of topic diversity and ensures the robustness of the topic modeling against meaningless word groupings.
@@ Appendix H Topic Distribution Visualization
Figure 32: This visualization represents a filtered subset of posts from the Danbooru2023 dataset, centered on the ’scenery’ tag. The network graph is an ego network (depth 1), which includes only nodes directly connected to the ’scenery’ tag. To refine the data and focus on meaningful associations, uncommon tags with fewer than 10 occurrences were excluded. The analysis, conducted using Gephi, focuses on nodes with a degree greater than 600 to highlight critical components. Nodes are color-coded by modularity class by Fast Unfolding Algorithm [10], revealing clusters of closely associated tags. Node size reflects Eigenvector Centrality[11], emphasizing highly connected and influential tags within their network.
Appendix I Discussion and Future Work
@@
Despite the promising performance demonstrated by TIPO, several areas merit further investigation:
Personalization and User Adaptation.
TIPO currently does not directly incorporate personalization methods (e.g. LoRA [30]). Incorporating such modules could further align generated outputs to individual user preferences, enabling more tailored creative assistance. Investigating joint training strategies that integrate personalized fine-tuning represents a valuable future direction.
Generalization to Uncommon Compositions.
While TIPO performs effectively on typical and broadly represented concepts, it may still struggle with highly unconventional combinations (e.g., “avocado chair”). Future research should systematically evaluate TIPO’s performance on rare or unusual prompts (but considered creative prompts), potentially exploring auxiliary training or adaptation methods to ensure broader compositional robustness.
Capacity and Backbone Scalability.
Given TIPO model’s architecture [66, 65], its capacity for handling highly complex or lengthy prompts could be inherently limited. Future work could investigate more extensive or specialized architectures enhance compositional and semantic capacity, ideally without compromising real-time inference efficiency, are crucial for practical deployment.

### v4

@@ G.5 Statistical Significance
G.6 Survey Response Examples
@@ G.7 Conclusion
Appendix H Ablation Study on TIPO
@@
H.1 Experimental Setup
Prompt Variants
@@ H.1 Experimental Setup
Inference Procedure
Prompts are fed into our image-generation pipeline under identical model settings (classifier free guidance, sampler, steps, etc.), using the v-parameterized variant of Illustrious v3.5(Park et al., 2024). We focus on how TIPO modifications alter the generation outcomes and whether they introduce additional computational overhead.
H.2 Evaluation Metrics
@@
H.3 Results & Discussion
Impact on Aesthetics
@@ H.3 Results & Discussion
Improvements in Corruption Score
@@
H.4 Speed Test and Overhead Analysis
Inference Speed
@@ H.4 Speed Test and Overhead Analysis
Memory Footprint
We also confirm that TIPO’s overhead in terms of VRAM usage is minimal (e.g., << 0.5 GB for TIPO-200M and << 1.5 GB for TIPO-500M) with the quantization supported by llama.cpp. The practical adoption of TIPO in pipelines has shown no critical memory concerns, which aligns with our measurements.
Training Costs
[…]
Although TIPO requires more total GPU-hours than PAE, it is trained on over 30 million prompts—two orders of magnitude larger than both Promptist and PAE. After normalization, TIPO achieves the lowest cost per 1k prompts, demonstrating strong scalability. It is also worth noting that Promptist and PAE rely on reinforcement learning with external T2I rollouts. Even for SD1.5, each rollout takes roughly five seconds, and the cost increases dramatically for larger models such as SDXL, SD3, or Flux. By contrast, TIPO requires no rollouts, so its training cost scales linearly with corpus size and remains independent of the target T2I model.
Table 13: Training cost comparison between TIPO and RL-based prompt optimization methods. GPU-hours per 1k prompts are normalized for fairness.
Method #Params #Prompts #GPUs GPU-h GPU-h /1k
[…]
H.5 Conclusion of Ablation
@@
Appendix I Topic Distribution Visualization
Latent Dirichlet Allocation (LDA) (Blei et al., 2003) is a generative probabilistic model for topic modeling (Jelodar et al., 2018), which assumes that each document is a mixture of topics, with each topic represented by a distribution over words. LDA uncovers hidden thematic structures by analyzing word co-occurrence patterns, while methods like TF-IDF and TextRank (Mihalcea and Tarau, 2004) enhance its ability to extract meaningful insights from large textual datasets. We implemented a multi-stage topic modeling and clustering methodology using LDA to extract varying numbers of topics (20, 30, 50, and 100) from the corpus. This approach focuses on identifying significant representative words while filtering out stop words and irrelevant terms to ensure meaningful topic classification.
We empirically assessed whether the resulting topics were sufficiently large and diverse by employing multi-level topic analysis. This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics (Stevens et al., 2012).
To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering (Zhao and Karypis, 2002). We grouped the initially extracted topics into five major clusters using k-means clustering. We evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) (Hartigan et al., 1979), shown in Table 14, 15, and 16. Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset. This two-tiered approach allows for a more nuanced analysis of topic diversity and ensures the robustness of the topic modeling against meaningless word groupings.
@@ Appendix I Topic Distribution Visualization
Figure 29: This visualization represents a filtered subset of posts from the Danbooru2023 dataset, centered on the ’scenery’ tag. The network graph is an ego network (depth 1), which includes only nodes directly connected to the ’scenery’ tag. To refine the data and focus on meaningful associations, uncommon tags with fewer than 10 occurrences were excluded. The analysis, conducted using Gephi, focuses on nodes with a degree greater than 600 to highlight critical components. Nodes are color-coded by modularity class by Fast Unfolding Algorithm (Blondel et al., 2008), revealing clusters of closely associated tags. Node size reflects Eigenvector Centrality(Bonacich, 1972), emphasizing highly connected and influential tags within their network.
Appendix J Discussion and Future Work
@@
Despite the promising performance demonstrated by TIPO, several limitations and future directions remain open for further study.
Distribution Dependence and OOD Generalization.
The optimization behavior of TIPO inherently depends on the distributional bias of the open text–image corpus used for training. As a result, it may exhibit unstable behavior on extremely rare or stylistically unconventional prompts. Besides, when applied to out-of-distribution (OOD) T2I models whose training data or prompting conventions deviate significantly from LAION-style distributions, aesthetic performance may degrade slightly. In addition, our current OOD evaluation uses GPT-4o-mini–generated prompts, which are high-quality but do not fully represent real-world user queries, thus limiting external validity. Addressing the generalization to unseen models and long-tail prompts remains an important direction for future work, which can be pursued from two complementary perspectives: improving the model’s generalization capacity via stronger backbones and domain-diverse fine-tuning, and enhancing evaluation through large-scale collection of authentic user prompts and cross-model benchmarking.
Stronger Backbone Initialization.
Our current implementation trains a mid-sized LLaMA variant from scratch. Future work could instead initialize from stronger open-source LLM backbones and fine-tune them on T2I corpora, potentially improving robustness on long-tail and domain-specific distributions. This could also mitigate failures on small or highly biased datasets by leveraging more general linguistic priors.
Model-Specific Adaptation and Style Variance.
For models that require structured or JSON-formatted prompts, TIPO can be combined with a lightweight adapter or fine-tuned on a small set of model-specific data. Extending TIPO with such adaptation modules could better accommodate systems whose prompt syntax diverges from mainstream diffusion models. In the longer term, TIPO can also serve as a backbone integrated with RL-based refinement for model-specific alignment.
Personalization and Style Preservation.
The current TIPO is a general-purpose optimizer and does not incorporate user-level or stylistic conditioning. Building on prior personalization techniques such as LoRA (Hu et al., 2022), future work could explore lightweight adapters or online learning mechanisms that track user preferences and maintain project-level stylistic consistency, enabling personalized and context-aware prompt optimization.
Image-Aware and Feedback-Driven Refinement.
TIPO currently operates purely in the text domain without utilizing generated images or user feedback. A promising extension is to incorporate vision–language models like Qwen3-VL (Bai et al., 2023) for image-aware refinement, allowing iterative refinement with optimized prompts, visual outcomes, and user instructions. However, such integration requires non-trivial data curation and training pipelines, which we leave for future work.
Scaling Behavior and Model Capacity.
From 200M to 500M parameters, TIPO continues to yield improvements in FDD and Aesthetic scores. Due to limited compute resources, we could not explore larger configurations to observe scaling saturation. A systematic study of TIPO’s scaling behavior and architectural variants would be a valuable direction for future research. Such analysis could reveal scaling laws unique to prompt optimizers and guide practical model sizing for future deployments.
Appendix K Disclosure of LLM Usage
We used GPT-5 only to polish writing by improving the readability and grammar correctness. No LLMs were used in the main contributions of this work, such as ideation, experiment design, or result analysis.